HETAV PATEL

Grade 11,

Ex 10.1 Class 6 Maths Question 6. Find the perimeter of each of the following shapes: (a) A triangle of sides 3 cm, 4 cm and 5 cm. (b) An equilateral triangle of side 9 cm. (c) An isosceles triangle with equal sides 8 cm each and third side 6 cm.

Ex 10.1 Class 6 Maths Question 6.
Find the perimeter of each of the following shapes:
(a) A triangle of sides 3 cm, 4 cm and 5 cm.
(b) An equilateral triangle of side 9 cm.
(c) An isosceles triangle with equal sides 8 cm each and third side 6 cm.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) We know that the perimeter of the given triangle = The sum of all sides of the triangle ∴ Perimeter of the triangle = 3 cm + 4 cm + 5 cm = 12 cm (b) We know that the perimeter of the given triangle = Sum of all the sides of the triangle = (9 + 9 + 9) = 27 cm (c) Perimeter of the given isosceles triangle = Sum of all the sides of the triangle = (8 + 8 + 6) cm = 22 cm